Would anyone mind helping me solve and provide an explanation for this problem? Thank You!

media%2F962%2F96287ab1-20cd-4392-b6c9-91

Would anyone mind helping me solve and provide an explanation for this problem? Thank You!
Answer #1

The charge of proton q= 1.60 * 10-19C

mass of proton m =1.60 * 10-27kg

The distance between two proton = r

The electric force between two protons Fe= (1/4*π*ε0)*(q*q/r2)

The gravitational force between two protons Fg= G*m*m / r2  

(A) the ratio between electric and gravitational force

Ratio = Fe/Fg= ((1/4*π*ε0)*(q*q/r2) /G*m*m / r2   ) = ((1/4*π*ε0)*(q*q) )/(G*m*m ) = 1.2 * 1036

(B) No, the attractive force of gravity is between two protons is very small in comparison with the electric force between them. It can not hold the nucleus together.

(C) Nucleus holds together because the nucleus is very small (1.2*10-14m in radius) in small distances strong inter-nucleus force works and this force is very strong then the electric force between two prints.

Because of this strong inter-nucleus force, It can hold the nucleus together.